# PAPER: Privacy-Preserving ResNet Models using Low-Degree Polynomial Approximations and Structural Optimizations on Leveled FHE

### v1

@@ 5 Experimental Results
5.1 Experimental Setup
We trained ResNet-18, ResNet-20, and ResNet-32 on CIFAR-10 and CIFAR-100 using PyTorch 2.4.1. Each experiment was repeated 10 times, and we report mean accuracy and private inference time with standard deviations. Private inference used a C++20 framework we developed, built on Microsoft SEAL 4.1 (SEAL, ) for leveled FHE operations and GMP 6.2.1 for multiprecision arithmetic. The framework includes a tool that automatically converts trained models for C++ inference. We use a machine with two AMD EPYC 64-core processors, 2 TB memory across 32 DDR4 DIMMs at 2933 MT/s, on Ubuntu 22.04.5 LTS with kernel 5.15.0-151-generic (x86_64) to run all experiments, including related work. For model training (see §2), we used hyperparameters $c=2$ and $\zeta=0.001$. An ablation study for both is provided in §G. We set $b=10$, aligned with PILLAR (Diaa et al., 2024), as it provides accurate polynomial approximation after quantization-aware fitting. For all clustering strategies, we used the $\ell_{2}$ norm as the distance metric, with k-means as the clustering algorithm (Lloyd, 1982). Details on FHE encryption parameters are in §H.1. Additional results, including peak memory, are reported in §H.3.
We compare our approach against non-interactive PPML baselines, MPCNN (Lee et al., 2022) and AutoFHE (Ao & Boddeti, 2024), in terms of accuracy and inference time. For fairness, we use their publicly available implementations (git, 2024a; 2022) with specified encryption parameters. Both baselines implement ResNet-20 on CIFAR-10 and ResNet-32 on CIFAR-10 and CIFAR-100, all with $N=2^{16}$. We also trained plaintext models using standard ReLU activations to serve as a reference for accuracy. In this section, we mainly report results for ResNet-20 and ResNet-32, since these allow direct comparison with baselines. Results for ResNet-18 are provided in §H.2.
5.2 Summary of Results
Accuracy of Polynomial Approximation. Training with the polynomial approximation described in §2 shows that, on average across CIFAR-10 and CIFAR-100, our degree-2 polynomial achieves accuracy only about 1.3% lower than ReLU, while PILLAR with degree-4 polynomial is about 5.4% lower. A degree-2 adaptation of PILLAR shows an average drop of 6.5%. These results demonstrate that our method can use low-degree polynomials effectively, while PILLAR does not reach comparable accuracy even with higher degrees. A detailed comparison is provided in §G.3.
@@ 5.2 Summary of Results
Figure 2: Accuracy vs. private inference time for ResNet-20/32 on CIFAR-10/100. Ensembles are evaluated with $M\in\{1,2,4\}$ ($M=1$ is a single model). We compare Standard ensembles (no clustering), Slice ensembles (slice-wise clustering), and Full Clustering. Colored squares represent centroid counts ($k=16,\dots,8192$), and the same color is used for the same centroid count in both slice and full clustering. Baselines include AutoFHE (numbers = bootstrapping layers) and MPCNN. The green dashed line shows plaintext ReLU accuracy.
@@
Ensemble Evaluation. We evaluate ensembles with $M=\{1,2,4\}$ models. Figure 2 (middle row) shows accuracy and inference time for two approaches: Standard-$M$ (ensembles without clustering) and Slice-$M$ (ensembles with slice clustering as in §4.3). Inference time does not increase with ensemble size since all models fit in one ciphertext, requiring the same number of homomorphic operations. For Standard-MM\!, accuracy consistently improves as $M$ increases. Slice-MM\!, however, does not always scale the same way. Since each centroid represents an $M$-dimensional space, higher $M$ can require more centroids to capture model parameters. Thus, Slice-4 can underperform Slice-2 at a fixed $k$, and Slice-2 needs more centroids to match Standard-2 accuracy. Despite this, Slice-$M$ provides notable advantages. Slice-2 mostly outperforms Standard-2 and reaches accuracy comparable to Standard-4. This suggests clustering acts as an implicit regularizer, mitigating overfitting.
@@
Comparison with Related Work. Figure 2 (bottom row) presents Pareto fronts of Slice Clustering in terms of accuracy and private inference time, compared with AutoFHE and MPCNN. For reference, we also report Standard-4 and plaintext ReLU accuracy. For ResNet-20, where our models operate with $N=2^{15}$, Slice Clustering achieves consistently faster inference than related work (up to $4\times$). For ResNet-32, where all approaches use $N=2^{16}$, our method is comparable in inference time to AutoFHE on CIFAR-10 and $2\times$ faster on CIFAR-100. Inference time does not change for our approach for different datasets, whereas AutoFHE slows down because it requires a high number of bootstrapping operations to maintain accuracy. The main factors affecting our inference time relative to related work are: The main factors affecting our inference time relative to related work are: (i) the polynomial activation (§2) and structural optimizations (§3) reduce both multiplicative depth and number of RNS moduli, allowing execution under LFHE without bootstrapping; (ii) our clustering technique (§4.2) further reduces the number of unique parameters, thereby limiting plaintext encodings per convolution; (iii) for ResNet-20 (and ResNet-18), we operate at a smaller polynomial degree, which directly accelerates homomorphic operations; (iv) our choice of data layout (§4.1) deliberately sacrifices some latency to better exploit ciphertext slots for accuracy, but above optimizations offset this overhead. In terms of accuracy, our approach consistently outperforms related work thanks to our training strategy (§2). In fact, it closely matches plaintext ReLU models, effectively eliminating the accuracy gap typically observed in polynomial approximations of ResNets.
@@ 6 Conclusion
Appendix A Leveled Fully Homomorphic Encryption
Homomorphic Encryption is a special type of encryption that allows meaningful computations to be performed directly on encrypted data. In this work, we employ the Cheon-Kim-Kim-Song (CKKS) (Cheon et al., 2017) encryption scheme, specifically the Residue Number System (RNS) variant (Cheon et al., 2018), which supports approximate arithmetic over encrypted real or complex numbers. CKKS is an FHE scheme based on the Ring-Learning With Errors (RLWE) problem (Lyubashevsky et al., 2013). The scheme operates across three domains: message $\mathcal{M}=\mathbb{C}$, plaintext $\mathcal{P}=R_{Q}=\mathbb{Z}_{Q}[x]/(x^{N}+1)$, and ciphertext $\mathcal{C}=R_{Q}\times R_{Q}$ spaces. The encoding function $encode(\cdot)=\mathcal{M}^{N/2}\mapsto\mathcal{P}$ maps a vector of $N/2$ complex numbers into a degree-$N$ polynomial in the plaintext space. The encryption function $encrypt(\cdot)=\mathcal{P}\mapsto\mathcal{C}$ transforms a plaintext into a ciphertext. In RNS-CKKS, the ciphertext modulus $Q$ is decomposed into $\mathcal{L}+1$ smaller primes $Q=\prod_{i=0}^{\mathcal{L}}{q_{i}}$, where $\mathcal{L}$ denotes the level. A polynomial in $R_{Q}$ is represented as $\mathcal{L}+1$ polynomials in $R_{q_{i}}\forall i\in[0,\mathcal{L}]\cap\mathbb{Z}$, which allows computations using native 64-bit integer arithmetic rather than costly arbitrary-precision operations.
CKKS supports three homomorphic operations: addition, multiplication, and rotation. Additions and multiplications act element-wise on the encoded vector, similar to SIMD vector operations. The rotation operation rotates the vector of $N/2$ encoded messages by a step $s\in\mathbb{Z}_{N/2}$. During encoding, coefficients are scaled by a factor $\Delta$, rounded to the nearest integers, and reduced modulo $Q$, making CKKS a fixed-point FHE scheme. Multiplying two encoded messages scales the result by $\Delta^{2}$, requiring a rescale operation to restore the scale to $\Delta$. Rescaling removes one RNS modulus, reducing the level by 1. It also accumulates approximation errors, since the RNS modulus being discarded and the scaling factor $\Delta$ are not equal. When the level reaches zero, bootstrapping can reset the ciphertext to a higher level, enabling unbounded computation. However, bootstrapping is computationally expensive and avoided when possible. FHE without bootstrapping is called Levelled Fully Homomorphic Encryption (LFHE). In LFHE, the encryption parameters $(N,Q)$ are chosen to provide the desired security and sufficient levels for a predefined arithmetic circuit with known multiplicative depth. To the best of our knowledge, this work is the first to apply LFHE to deep CNN models like ResNet-20 and ResNet-32.
Appendix B Challenges of Using Polynomial Activation
ReLU, defined as $ReLU(x)=\max(0,x)$, is one of the most widely used activation functions in deep learning due to its simplicity and computational efficiency. Its piecewise linear structure introduces the non-linearity necessary for neural networks to capture complex patterns in data. While ReLU is straightforward to implement in plaintext settings, it poses significant challenges in FHE environments due to its reliance on conditional branching. FHE can evaluate only polynomial functions as it supports only additions and multiplications, thus, any FHE representation of ReLU must be a polynomial approximation. Evidently, the higher degree the polynomial, the more precise the approximation. However, the multiplicative depth of the activation functions grows logarithmically to the degree of the polynomial used for approximation. Precisely, the multiplicative depth can be computed as $\delta=\lceil\log_{2}{d}+1\rceil$, where $d$ denotes the polynomial degree and the +1+1 refers to the coefficient multiplication. In the literature, the lowest multiplicative depth for a polynomial approximation of ReLU is achieved by PILLAR (Diaa et al., 2024), which uses a degree-4 polynomial of the form $\sum_{i=0}^{4}{c_{i}x^{i}}$ with $\delta=3$. The theoretical minimum multiplicative depth for an activation function is one, achieved with a polynomial of the form $x^{2}+c_{1}x+c_{0}$. Polynomials like this have been effectively used in shallow CNNs like CryptoNets (Gilad-Bachrach et al., 2016) and LoLa (Brutzkus et al., 2019), but failed to work for deeper models due to the escaping activation problem (Garimella et al., 2021). In addition, polynomial approximations are inherently limited to bounded intervals of the input domain. When substituted directly for ReLU during model training, these approximations often lead to a significant drop in model accuracy (Garimella et al., 2021; Hussain et al., 2021; Diaa et al., 2024). This section analyzes two primary causes of this degradation: escaping activations, where intermediate layer outputs fall outside the polynomial’s approximation interval and coefficient truncation, which results from representing polynomial coefficients in fixed-point format to ensure compatibility with FHE arithmetic.
B.1 Escaping Activation Problem
Let $p_{d}(x)=\sum_{k=0}^{d}a_{k}x^{k}$ denote a degree-$d$ polynomial approximation of the ReLU function, where each coefficient $a_{k}\in\mathbb{R}$ for all $k\in\{0,\dots,d\}$. These coefficients are typically obtained by minimizing the least-squares error between the polynomial and the ReLU function over a finite set of real-valued sample points $\{x_{i}\}_{i=1}^{N}\subset[-c,c]$, for some parameter $c>0$:
$\min_{a_{0},\ldots,a_{d}}\;\sum_{i=1}^{N}\left(ReLU(x_{i})\;-\;p_{d}(x_{i})\right)^{2}$
The resulting polynomial $p_{d}(x)$ provides a close approximation to ReLU within the interval [c,c][-c,c]. However, outside the interval, the polynomial behavior differs significantly from that of ReLU. ReLU exhibits piecewise linear growth:
In contrast, the growth of $p_{d}(x)$ for $d\geq 2$ is polynomial, dominated asymptotically by the highest-degree term:
$p_{d}(x)\;=\;a_{d}x^{d}\;+\;a_{d-1}x^{d-1}\;+\;\cdots\;+\;a_{1}x\;+\;a_{0}\;=\;\mathcal{O}(x^{d})$
This fundamental mismatch in growth rates gives rise to the problem of escaping activations, first identified by Garimella et al. (2021). As inputs propagate through the layers of a neural network, the intermediate values passed into the polynomial activation can escape the intended interval [c,c][-c,c], entering regions where $p_{d}(x)$ no longer approximates ReLU accurately. As a consequence, the network can produce excessively large activations, which in turn cause exploding weights and rapid degradation of model performance unless specific modifications are made to the training procedure to contain them.
B.2 Coefficient Truncation
In privacy-preserving inference based on FHE, all computations are carried out using fixed-point arithmetic with limited precision. This constraint restricts the range and resolution of values that can be accurately represented, affecting both the domain of activation function inputs and the magnitudes of polynomial approximation coefficients. The polynomial coefficients $\{a_{k}\}_{k=0}^{d}$ are typically derived via floating-point least-squares fitting. To enable fixed-point evaluation under FHE, these coefficients are quantized using:
$\tilde{a}_{k}\;=\;\frac{\lfloor a_{k}\cdot 2^{b}\rceil}{2^{b}}$
where $b\in\mathbb{N}$ represents the number of fractional bits in the fixed-point format and $\lfloor\cdot\rceil$ denotes rounding to the nearest integer. The least-squares fitting procedure often produces small-magnitude coefficients, particularly for higher-degree monomials. If any coefficient satisfies $\left|a_{k}\right|<2^{-(b+1)}$, the quantized value $\tilde{a}_{k}$ becomes zero. This effectively discards the corresponding monomial $x^{k}$ from the approximation. This phenomenon, referred to as coefficient truncation (Diaa et al., 2024), changes the shape of the polynomial and can cause significant deviation from the intended ReLU behavior, even within the designated approximation interval [c,c][-c,c].
Appendix C Why the Penalty Function Works?

### v2

@@ 6 Experimental Results
6.1 Evaluation of Training Methodology
Training Setup. We evaluate our polynomial approximation strategy on the CIFAR-10, CIFAR-100, and Tiny-ImageNet (referred to as Tiny) datasets, which are widely adopted in related work on privacy-preserving neural network inference. Our experiments use ResNet-18, ResNet-20, and ResNet-32 architectures, and additionally include VGG-16 to assess generality beyond residual networks. Each experiment is repeated 10 times with different random seeds, and we report mean classification accuracy along with standard deviations. For model training (see §3), we set the clipping parameter to $c=2$ and the penalty strength to $\zeta=0.001$; the impact of both hyperparameters is analyzed subsequently. We fix the fixed-point precision to $b=10$, following PILLAR [11], as it provides accurate polynomial approximation under quantization-aware fitting. For all clustering strategies, we use the $\ell_{2}$ norm as the distance metric and employ k-means clustering [22].
Impact of Clipping Parameter ($c$). We evaluate the effect of the clipping interval [c,c][-c,c] on model performance. The choice of $c$ determines the trade-off between preserving the activation distribution and truncating extreme values. If $c$ is too small, informative activations are excessively clipped, reducing representational capacity; conversely, if $c$ is too large, the clipping mechanism provides little stabilization, allowing unstable activations to propagate. Results are reported for ResNet-18 on CIFAR-10, which we present without loss of generality. Figure 7(a) shows the distribution of pre-activation inputs to ReLU across the network. The values are largely concentrated within the interval [2,2][-2,2], with only a small fraction outside this range, supporting the choice of restricting pre-activations to [2,2][-2,2]. Figure 7(b) shows the classification accuracy obtained under different clipping intervals. Among the tested ranges, [2,2][-2,2] consistently achieves the highest mean accuracy across runs. Smaller intervals degraded performance by discarding a substantial portion of activations, while larger intervals reduce stability and result in lower accuracy. This trend is consistent across other dataset–architecture combinations. Based on these observations, we adopt [2,2][-2,2] as the clipping range in all experiments.
Impact of Penalty Strength ($\zeta$). We analyze the impact of the penalty strength $\zeta$ on model performance. The parameter $\zeta$ controls the magnitude of the regularization term that penalizes pre-activations outside the target interval [c,c][-c,c]. Figure 7(c) reports results ResNet-18 on CIFAR-10. The highest mean accuracy is obtained for $\zeta=10^{-3}$. Smaller values fail to provide sufficient regularization and lead to degraded accuracy, whereas larger values reduce performance by overly constraining the model. This trend is consistent across other dataset–architecture combinations. Based on these observations, we set $\zeta=10^{-3}$ in all experiments.
Figure 7: Effects of activation clipping and regularization for ResNet-18 on CIFAR-10. (a) Pre-Activation value distribution across all layers, showing concentration within [-2, 2]. (b) Accuracy across clipping ranges, with [-2, 2] performing best. (c) Accuracy versus penalty strength $\zeta$, with best performance at $\zeta=10^{-3}$; smaller values under-regularize the model, while larger values overly constrain it and reduce accuracy.
Comparison to Related Work. We compare the training methodology of our approach against PILLAR, which represents the lowest-degree polynomial-approximation method previously shown to train deep neural networks reliably. We evaluate PILLAR using its open-source implementation [3]. PILLAR was originally designed in the context of interactive PPML, which differs from our non-interactive setting, and its repository does not provide any implementation for private inference. For these reasons, we restrict the comparison to classification accuracy and do not evaluate inference time. Table 2 reports classification accuracies (mean $\pm$ standard deviation) across all evaluated datasets and architectures using ReLU, PAPER (our method), and two versions of PILLAR. The original PILLAR method employs a degree-4 polynomial activation. For a fair comparison, we additionally adapt this implementation to use a degree-2 polynomial, allowing a direct comparison with PAPER at the same polynomial degree while keeping the original training procedure unchanged. As shown in Table 2, when averaged across all datasets and architectures, the adapted degree-2 variant of PILLAR exhibits a substantial degradation in accuracy, with an average drop of 7.99% relative to ReLU and notably high variance, particularly on more challenging datasets such as CIFAR-100 and Tiny-ImageNet. The original degree-4 variant of PILLAR is more stable but still underperforms ReLU, with an average accuracy gap of 6.54%. In contrast, PAPER with degree-2 polynomial activation remains much closer to ReLU, incurring an average accuracy drop of only 1.84% across all datasets and architectures. Importantly, PAPER demonstrates consistently stable training behavior across all evaluated settings, avoiding the large run-to-run fluctuations seen in both variants of PILLAR. These results demonstrate that while PILLAR either degrades significantly when reduced to a degree-2 configuration or remains less accurate in its intended degree-4 configuration, PAPER enables accurate and stable training with quadratic polynomial activations. Overall, these results indicate that the accuracy gap typically associated with low-degree polynomial approximations arises primarily from limitations in training methodology, and can be substantially mitigated through the regularization techniques introduced in PAPER.
Table 2: Classification accuracy (%, mean $\pm$ standard deviation) on CIFAR and Tiny-ImageNet datasets for VGG and ResNet models using ReLU, PAPER (degree-2), and two versions of PILLAR.
Method VGG-16 ResNet-18 ResNet-20 ResNet-32
CIFAR-10
ReLU 92.57 $\pm$ 0.36 94.57 $\pm$ 0.23 94.84 $\pm$ 0.25 95.23 $\pm$ 0.30
PAPER 90.41 $\pm$ 0.46 93.73 $\pm$ 0.27 94.06 $\pm$ 0.26 93.86 $\pm$ 0.24
PILLAR ($d=2$) 88.32 $\pm$ 0.19 91.78 $\pm$ 0.13 90.43 $\pm$ 0.40 90.40 $\pm$ 1.73
PILLAR ($d=4$) 90.76 $\pm$ 0.18 93.26 $\pm$ 0.15 90.98 $\pm$ 0.29 92.24 $\pm$ 0.27
Method ResNet-18 ResNet-20 ResNet-32 ResNet-32
CIFAR-100 Tiny
ReLU 76.22 $\pm$ 0.44 74.96 $\pm$ 0.36 76.58 $\pm$ 0.40 61.37 $\pm$ 0.36
PAPER 74.96 $\pm$ 0.51 73.89 $\pm$ 0.37 73.93 $\pm$ 0.53 56.78 $\pm$ 0.42
PILLAR ($d=2$) 71.92 $\pm$ 0.28 62.30 $\pm$ 3.96 66.38 $\pm$ 1.78 40.85 $\pm$ 0.68
PILLAR ($d=4$) 74.81 $\pm$ 0.22 62.55 $\pm$ 7.51 66.10 $\pm$ 1.12 43.29 $\pm$ 0.83
Note: The original PILLAR method uses a degree-4 polynomial [11]; the degree-2 variant is our adaptation of its open-source code for comparison.
6.2 Evaluation of Private Inference
Private Inference Setup. Private inference uses an in-house C++20 framework, built on Microsoft SEAL 4.1 [32] for leveled FHE operations and GMP 6.2.1 for multiprecision arithmetic. The framework includes a tool to automatically convert trained models for C++ inference. All experiments, including related work, were run on a machine with two AMD EPYC 64-core processors, 2 TB memory across 32 DDR4 DIMMs at 2933 MT/s, on Ubuntu 22.04.5 LTS. For PAPER, we set $N=2^{16}$ for ResNet-32 and $N=2^{15}$ for the other models, as these are the smallest polynomial degrees enabling correct inference with $\approx$128-bit security. Details on FHE encryption parameters are in §C. We compare our approach against state-of-the-art non-interactive PPML baselines, MPCNN [20] and AutoFHE [5], in terms of accuracy and inference time, using their publicly available implementations [2, 1] as is. Both baselines implement ResNet-20 on CIFAR-10 and ResNet-32 on CIFAR-10 and CIFAR-100, all with $N=2^{16}$. We also train plaintext models using standard ReLU activations to serve as reference for accuracy. Figure 8 summarizes private inference time versus accuracy across several CNN models and datasets, comparing clustering methods, ensembles, and baselines.
Figure 8: Accuracy vs. private inference time for VGG and ResNet models on CIFAR and Tiny-ImageNet datasets. Ensembles are evaluated with $M\in\{1,2,4\}$ ($M=1$ is a single model). We compare Standard ensembles (no clustering), Slice ensembles (Slice Clustering), and Full Clustering. Colored squares represent centroid counts ($k=16,\dots,8192$), and the same color is used for the same centroid count in both slice and full clustering. Baselines include AutoFHE (numbers = bootstrapping layers) and MPCNN. The green line shows plaintext ReLU accuracy.
@@ 6.2 Evaluation of Private Inference
Ensemble Evaluation. In Figure 8 (rows 2 and 5), we evaluate ensembles with $M=\{1,2,4\}$ models. The figure depicts accuracy and inference time for two approaches: Standard-$M$ (ensembles without clustering) and Slice-$M$ (ensembles with slice clustering as in §5.3). Inference time does not increase with ensemble size since all models fit in one ciphertext, requiring the same number of homomorphic operations. For Standard-MM\!, accuracy consistently improves as $M$ increases. Slice-MM\!, however, does not always scale the same way. Since each centroid represents an $M$-dimensional space, higher $M$ can require more centroids to capture model parameters. Thus, Slice-4 can underperform Slice-2 at a fixed $k$, and Slice-2 needs more centroids to match Standard-2 accuracy than either method with $M=1$. Despite this, Slice-$M$ is advantageous as it produces models with higher accuracy. For instance, Slice-2 outperforms Standard-2 and reaches accuracy comparable to Standard-4. This suggests clustering acts as an implicit regularizer that mitigates overfitting.
Accuracy of Polynomial Approximation. Figure 8 (rows 3 and 6) presents Pareto fronts for Slice Clustering in terms of accuracy and private inference time, and compares them against AutoFHE and MPCNN. For reference, the figure also reports Standard-4 and plaintext ReLU accuracy. Beyond the gains obtained from our training methodology, Slice Ensemble Clustering5.3) further improves accuracy by ensembling multiple models within a single ciphertext. For the simpler CIFAR-10 dataset, Slice Clustering achieves accuracy ranging from 1.1% below to 0.3% above ReLU. For the more complex CIFAR-100 and Tiny-ImageNet datasets, our approach consistently surpasses ReLU by 0.7–2.6%. These results demonstrate that our method enables the effective use of low-degree polynomial activations without sacrificing accuracy.
@@
Comparison to Related Work. For ResNet-20, where our models operate with $N=2^{15}$, Slice Clustering achieves consistently faster inference than related work (up to $4\times$). For ResNet-32, where all approaches use $N=2^{16}$, our method is comparable in inference time to AutoFHE on CIFAR-10 and $2\times$ faster on CIFAR-100. Inference time does not change for our approach for different datasets, whereas AutoFHE slows down because it requires a high number of bootstrapping operations to maintain accuracy. The main factors affecting our inference time relative to related work are: (i) the polynomial activation (§3) and structural optimizations (§4) reduce both multiplicative depth and number of RNS moduli, allowing execution under LFHE without bootstrapping; (ii) our clustering technique (§5.2) further reduces the number of unique parameters, thereby limiting plaintext encodings per convolution; (iii) for ResNet-20, we operate at a smaller polynomial degree, which directly accelerates homomorphic operations; (iv) our choice of data layout (§5.1) deliberately sacrifices some latency to better exploit ciphertext slots for accuracy, but the above optimizations offset this overhead. In terms of accuracy, our approach consistently outperforms related work thanks to our training strategy (§3). In fact, it closely matches plaintext ReLU models, effectively eliminating the accuracy gap typically observed in polynomial approximations of CNNs.
Scalability. The training methodology and optimization techniques exhibit scalability along multiple dimensions: they (i) are applicable to different CNN architectures, as demonstrated with VGG and ResNet models; (ii) extend to deeper networks, allowing more layers before bootstrapping, as shown with ResNet-32, whereas prior LFHE works were limited to AlexNet and VGG-11 [12]; (iii) support datasets with both a small (CIFAR-10) and a large number of classes (CIFAR-100); and (iv) accommodate higher-resolution images (Tiny-ImageNet) with no inference slowdown provided the number of ciphertexts for data encoding remains unchanged.
@@ 7 Conclusion
Ethical Considerations
Our work focuses on improving the computational efficiency and accuracy of privacy-preserving machine learning using leveled FHE. All experiments use publicly available benchmark datasets such as CIFAR-10, CIFAR-100, and Tiny-ImageNet. The research does not involve human subjects, personal data, or interactions with live systems. The primary stakeholders are end users who need strong protection for sensitive data, organizations that provide encrypted inference services, and the research community. Our techniques reduce reliance on costly bootstrapping and allow deeper neural networks to run efficiently on encrypted inputs. This provides a clear societal benefit by making practical privacy-preserving machine learning more accessible. The method operates strictly within the security guarantees of the underlying cryptographic scheme and does not introduce new vulnerabilities or mechanisms that could enable data misuse or surveillance. It focuses solely on improving the performance of established secure computation methods. All implementation details and model configurations are included to support reproducibility. Because the research does not involve human subjects, sensitive information, or activities that require institutional oversight, no IRB review is required.
Appendix A Why the Penalty Function Works?